C. Find the probabilities on a standard normal curve.

1. P(-3.0 ˂ z < -1.68)

2. P(-2.76 < z < 2)

3. P(1.53 < z < 2)


1
Expert's answer
2022-04-06T09:18:11-0400

"1.\ P(-3.0<z<-1.68)=\\=P(z<-1.68)-P(z<-3.0)="

"=0.04648-0.00135=0.04513" ​(from z-table)


"2.\ P(-2.76<z<2)=\\=P(z<2)-P(z<-2.76)="

"=0.97725-0.00289=0.97436" ​(from z-table)


"3.\ P(1.53<z<2)=\\=P(z<2)-P(z<1.53)="

"=0.97725-0.93699=0.04026" ​(from z-table)